OpenStudy (anonymous):

simplify: (5/4x^2y) - (y/14xz)

OpenStudy (anonymous):

\[\frac{5}{4x^2y} - \frac{y}{14xz}=\frac{35z}{28x^2yz} - \frac{2xy^2}{28x^2yz}\]\[=\frac{35z-2xy^2}{28x^2yz}\]